# Boosting Prediction with Data Missing Not at Random

### v2

@@ 1 Introduction
Boosting, a useful machine learning method, transforms weak learners into strong learners through iterative processes (Schapire and Freund,, 2012). The AdaBoost algorithm (Freund and Schapire,, 1997) is widely recognized as the first practically feasible boosting algorithm, regarded as the best off-the-shelf classifier (Breiman,, 1998). Breiman, (1998, 1999) demonstrated that the AdaBoost algorithm can be interpreted as a steepest descent algorithm in a function space spanned by weak learners. Friedman et al., (2000) and Friedman, (2001) developed a general statistical framework, providing a direct interpretation of boosting as a function estimation method. They also extended boosting from binary classification to regression and multiclass classification. Bühlmann and Yu, (2003) introduced a computationally simple boosting algorithm for regression and classification problems using the $L_{2}$ loss function.
These boosting methods typically require access to complete data without missing values, which is, however, often not true in practice. Recognizing this limitation, Hothorn et al., (2006), Barnwal et al., (2022), Chen and Yi, (2024), Bian et al., 2025a , among others, extended boosting algorithms to accommodate censored response variables. In this paper, we focus on extending boosting algorithms to handle incomplete data scenarios where the response variable is subject to missingness.
When data are missing completely at random (MCAR), conventional boosting algorithms can be directly applied to the complete observations, yielding valid results as they still constitute a random subsample. However, directly applying boosting procedures to data missing not at random (MNAR) poses a significant challenge, as it can lead to biased results (Bian et al., 2025b, ). This issue is further complicated by the inherent non-identifiability problems associated with MNAR data, as discussed by Wang et al., (2014), Miao and Tchetgen Tchetgen, (2016), Sun et al., (2018), and Morikawa and Kim, (2021), among others.
In this paper, we incorporate two strategies to adjust the loss function: the Buckley-James-type and the inverse propensity weight adjustments, with the missingness effects accounted for. These adjustments allow the flexible application of boosting methods to incomplete data with MNAR through a functional gradient descent algorithm. However, the usual optimization procedures are complicated by the involvement of unknown functions. To address this difficulty, we describe semiparametric optimal estimation approaches that provide consistent estimators for these unknown functions. We rigorously establish the theoretical properties for the resultant estimators. In addition, we adapt conditions considered by Morikawa and Kim, (2021) to ensure model identifiability.
@@ 2.1 Objective and Data
Let $Y$ denote the continuous response variable, and let $X$ denote the $p$-dimensional random vector of covariates, where $Y\in\mathcal{Y}$, $X\in\mathcal{X}$, and $\mathcal{Y}$ and $\mathcal{X}$ are the sample spaces for $Y$ and $X$, respectively. The goal is to find a function of $X$, say $f(X)$, that can effectively predict $Y$. Let =$\mathcal{F}=\{f:\mathcal{X}\to\mathcal{Y}\ |\ f\text{ is continuous and bounded in }\ell_{\infty}\}$ denote the set of real valued functions satisfying condition (B9) in the Supplementary Materials. For $f\in\mathcal{F}$, let $L:\mathcal{Y}\times\mathcal{Y}\xrightarrow{\ }\mathbb{R}$ denote the loss function that describes the discrepancy of using $f(X)$ to predict $Y$, which is assumed to be differentiable (almost everywhere) and convex with respect to the second argument, as specified in condition (B3) of the Supplementary Materials.
@@ 2.2 Conventional Boosting Prediction Procedure
Finding $\hat{f}_{\text{full}}$ in (3) can be achieved by employing the boosting method (Freund and Schapire,, 1997). This method, also known as the functional gradient descent algorithm (e.g., Bühlmann and Yu,, 2003; Bühlmann and Hothorn,, 2007; Schapire and Freund,, 2012), minimizes the empirical risk through steepest gradient descent in a function space to iteratively improve the estimate of $\hat{f}_{\text{full}}$ (Friedman,, 2001), where the function space is spanned by a class of constrained continuous functions, denoted by $\mathcal{C}$. At iteration $m$, given the current estimate $f^{(m)}(\cdot)$, the next estimate $f^{(m+1)}$ of $\hat{f}_{\text{full}}$ is updated by adding an increment term, $\hat{\alpha}^{(m+1)}\hat{h}^{(m+1)}(\cdot)$, where $\hat{h}^{(m+1)}(\cdot)\in\mathcal{C}$, and $\hat{\alpha}^{(m+1)}$ is a learning rate.
@@
The iteration procedure terminates when a specified stopping criterion is met, say after $\tilde{m}$ iterations. The final estimator of $f^{*}$ is then given by $\hat{f}_{\text{full}}(\cdot)\triangleq f^{(\tilde{m})}(\cdot)$. A common stopping criterion evaluates the difference in the values of $L(\cdot,\cdot)$ between successive estimates, $f^{(m)}(\cdot)$ and $f^{(m-1)}(\cdot)$. The iteration stops when this difference, measured in a certain norm (e.g., $L_{1}$ or $L_{2}$), falls below a prespecified threshold value (e.g., $10^{-6}$).
@@ 3.1 Missing not at Random and Identifiablity
Let $f(y|X=x)$ denote the conditional probability density of $Y$ given $X$, and let $\pi(y,x)\triangleq\Pr(R=1|Y=y,X=x)$ denote the propensity of observing the response, i.e., the conditional probability of observing $Y$, given $Y$ and $X$. If $\pi(y,x)$ does not depend on $Y$, the resulting missing data mechanism is called missing at random (MAR). When the missing mechanism is missing not at random (MNAR) or nonignorable, the propensity $\pi(y,x)$ depends on $Y$, as well as $X$, irrespective of whether $Y$ is missing or observed. Under MNAR, non-identifiability is often a concern (Robins and Ritov,, 1997). When both $f(y|X=x)$ and $\pi(y,x)$ are left fully unspecified, the joint distribution of $Y$ and $R$, given $X$, becomes non-identifiable (Robins and Ritov,, 1997).
To address non-identifiability issues, certain assumptions can be imposed. For example, Wang et al., (2014) assumed the existence of a nonresponse instrumental variable (aka a shadow variable) (Miao and Tchetgen Tchetgen,, 2016), which is a component of the covariate vector $X$ that is associated with the response $Y$ but is conditional independent of the missingness indicator $R$, given $Y$ and other components of $X$. Alternatively, Sun et al., (2018) assumed the existence of another version of instrumental variables: a subset of the covariate vector $X$ that is independent of the response $Y$ but conditionally dependent on $R$, given $Y$ and other components of $X$. While utilizing instrumental variables can mitigate non-identifiability issues, identifying an appropriate instrumental variable can be difficult in applications. Most importantly, those conditions are not testable solely based on the observed data (Morikawa and Kim,, 2021).
Instead of relying on the existence of instrumental variables, Morikawa and Kim, (2021) proposed a set of identification conditions, listed as (A1) - (A4) in the Supplementary Materials. Those conditions are testable using observed data. Even if the model does not satisfy those conditions, a doubly-normalized exponential transformation (Morikawa and Kim,, 2021) can be applied to artificially make the model identifiable, albeit at the cost of sacrificing the estimator consistency. In our development here, we adopt the identification conditions of Morikawa and Kim, (2021) to emphasize our key ideas.
Imposing parametric assumptions on both $f(y|X=x)$ and $\pi(y,x)$ can help address the non-identifiability issue, but the results are often sensitive to model misspecification (Kenward,, 1998). As a remedy, an intermediate approach may be considered, where one function is modeled parametrically while leaving the other unspecified; this is the strategy we adopt in the following development.
@@ 3.2 Adjusting Loss Functions with MNAR Data
The boosting prediction procedure described in Section 2.2 cannot be applied directly to $\mathcal{O}_{\text{missing}}$, as the response $Y_{i}$ is not observed for every subject in the study. To address this, we construct a new loss function, denoted $L^{*}(y_{i},f(x_{i}),r_{i})$, using the observed data in the sample $\mathcal{O}_{\text{missing}}$. Following the idea of Chen and Yi, (2024), we construct an adjusted loss function $L^{*}(\cdot,\cdot,\cdot)$ that maintains the same risk function as the original loss function $L(\cdot,\cdot)$, i.e., $E\{L^{*}(Y_{i},f(X_{i}),R_{i})\}=E\{L(Y_{i},f(X_{i}))\}$. Therefore, minimizing the risk function of the adjusted loss function $E\{L^{*}(Y_{i},$ $f(X_{i}),R_{i})\}$ is equivalent to minimizing the risk function $R(f)$ defined in (1), as if all $Y_{i}$ were observed.
Following Hothorn et al., (2006), one way to adjust the loss function is through the inverse propensity weight (IPW) scheme:
@@
To maximize the use of all available measurements, we follow the idea of Chen and Yi, (2024) to construct a Buckley-James (BJ)-type adjusted loss function:
@@
[…]
Determining $\Psi_{0}(x)$ requires $f(y|X=x,R=0)$, which is not available since the outcome is not observed for this subpopulation. To get around this, as in Kim and Yu, (2011), we use Bayes’ rule to express
@@ 4.1 Construction of Consistent Estimators
Since some $y_{i}$ values are missing (i.e., when $r_{i}=0$), directly setting (8) to 0 to solve for $\gamma$ is not feasible. To address this, Morikawa and Kim, (2021) proposed an alternative estimation method. Let
@@
Let $S_{0}(Y,X;\gamma)$ denote (9) with $r$ set to 0. Let $\hat{E}^{*}\{S_{0}(Y,X;\gamma)|X=x\}$ represent an estimate of (10), with $E(\cdot|X=x,R=1)$ estimated using the conditional probability density function $f(y|X=x,R=1)$.
@@
$\sum^{n}_{i=1}\left[\left\{1-\frac{r_{i}}{\pi(y_{i},x_{i};\gamma)}\right\}\hat{E}^{*}\{S_{0}(Y_{i},X_{i};\gamma)|X_{i}=x_{i}\}\right]=0$ (11)
@@
With $\hat{\beta}_{\text{p}}$, (10) can be estimated accordingly. If the conditional expectations in (10) can be derived in closed-form, then $\hat{E}^{*}\{S_{0}(Y,X;\gamma)|X=x\}$ in (11) is estimated as:
@@
where the conditional expectations are evaluated with respect to $f(y|X=x,R=1;\hat{\beta}_{\text{p}})$. Otherwise, $\hat{E}^{*}\{S_{0}(Y,X;\gamma)|X=x\}$ can be approximated using the fractional weights approach (Kim,, 2011) by
@@
Under identification conditions (A1) - (A4) and regularity conditions (C1) - (C6), listed in the Supplementary Materials, a consistent estimator of $\gamma$, denoted $\hat{\gamma}_{\text{p}}$, can be obtained by solving (11), with $\hat{E}^{*}\{S_{0}(Y,X;\gamma)|X=x\}$ replaced by (12) or (13). Hence, $\hat{\pi}_{\text{p}}(y,x)\triangleq\pi(y,x;\hat{\gamma}_{\text{p}})$ provides a consistent estimate for $\pi(y,x)$. While the performance of $\hat{f}_{\text{p}}(y|X=x,R=1)$ may be sensitive to the assumed parametric form $f(y|X=x,R=1;\beta)$, Morikawa and Kim, (2021) justified that $\hat{\gamma}_{\text{p}}$ remains consistent even if $f(y|X=x,R=1;\beta)$ is misspecified. Moreover, when $f(y|X=x,R=1;\beta)$ is correctly specified, $\hat{\gamma}_{\text{p}}$ attains the semiparametric efficiency bound.
@@
Under identification conditions (A1) - (A4) and regularity conditions (C1) - (C3) and (C7) - (C12), stated in the Supplementary Materials, a consistent estimator of $\gamma$, denoted $\hat{\gamma}_{\text{np}}$, which attains the semiparametric efficiency bound, can be obtained by solving (11) with $\hat{E}^{*}\{S_{0}(Y_{i},X_{i};\gamma)|X_{i}=x_{i}\}$ replaced by (14). Therefore, $\hat{\pi}_{\text{np}}(y,x)\triangleq\pi(y,x;\hat{\gamma}_{\text{np}})$ is a consistent estimator for $\pi(y,x)$.
@@ 4.3 Specification of the class $\mathcal{C}$
The boosting prediction procedure for MNAR data described in Section 4.2 depends on the specification of the class $\mathcal{C}$. Similar to Chen and Yi, (2024), we consider the regression spline method to characterize the functions in $\mathcal{C}$. Instead of using the truncated power basis functions, which are known to be numerically unstable in practice (Hastie et al.,, 2009), we employ more computationally stable B-spline basis functions (de Boor,, 2001).
@@
with $h_{0}$ denoting the intercept and $h_{k}(x_{k})=\sum^{T}_{t=1}\xi_{kt}b_{kt}(x_{k})$ for $k=1,\ldots,p$, where $x=(x_{1},\ldots,x_{p})^{\top}$; $b_{k}(x_{k})=(b_{k1}(x_{k}),\ldots,$ $b_{kT}(x_{k}))^{\top}$ is the vector of the B-spline basis functions of order $M$, with $T-M+1$ interior knots at suitably chosen quantiles of $X_{k}$ and $M\geq 2$ being an integer; and $\xi_{k}=(\xi_{k1},\ldots,\xi_{kT})^{\top}$ is the unknown parameter vector.
@@ 5 Theoretical Results
While the convergence of Algorithm 1 is guaranteed by Theorem 1, in applications, Algorithm 1 stops at a finite number to avoid excessively running an unnecessarily large number of iterations. Similar to Chen and Yi, (2024), we may use (18), in combination with (17), to decide an upper bound for the stopping iteration number $\tilde{m}$ in Algorithm 1:
@@
where $\eta$ is a given threshold in Algorithm 1. The upper bound of (19) shows that the number of iterations may vary with the choice of an initial function $f^{(0)}$ and the required accuracy $\eta$, as noted by Boyd and Vandenberghe, (2004) and Chen and Yi, (2024).
@@ 6.1 Simulation Design and Data Generation
We consider two settings, as in Bian et al., 2025b . with different dimensions $p$ for covariates $X_{i}$: Setting 1 has $p=2$, and Setting 2 includes a larger number of covariates with $p=9$.
@@
For Setting $j$ with $j=1,2$, given the covariates $X_{i}$, we generate the missing data indicator $R_{i}$ from Bernoulli($\tilde{\pi}_{j}(X_{i})$), with
@@
Condition (A4) in the Supplementary Materials implies that setting $\beta_{1,3}\neq 0$ ensures the model to be identifiable. Noted by Morikawa and Kim, (2021) through simulation studies, almost half of the simulations failed to converge if $\beta_{1,3}$ is taken as 0. Consequently, here we only consider settings with a nonzero $\beta_{1,3}$, set as $\beta_{1,3}=0.8$.

### v3

@@ 1 Introduction
Boosting, a useful machine learning method, transforms weak learners into strong learners through iterative processes (Schapire and Freund, 2012). The AdaBoost algorithm (Freund and Schapire, 1997) is widely recognized as the first practically feasible boosting algorithm, regarded as the best off-the-shelf classifier (Breiman, 1998). Breiman (1998, 1999) demonstrated that the AdaBoost algorithm can be interpreted as a steepest descent algorithm in a function space spanned by weak learners. Friedman et al. (2000) and Friedman (2001) developed a general statistical framework, providing a direct interpretation of boosting as a function estimation method. They also extended boosting from binary classification to regression and multiclass classification. Bühlmann and Yu (2003) introduced a computationally simple boosting algorithm for regression and classification problems using the $L_{2}$ loss function.
These boosting methods typically require access to complete data without missing values, which is, however, often not true in practice. Recognizing this limitation, Hothorn et al. (2006), Barnwal et al. (2022), Chen and Yi (2024), and Bian et al. (2025a), among others, extended boosting algorithms to accommodate censored response variables. In this paper, we focus on extending boosting algorithms to handle incomplete data scenarios where the response variable is subject to missingness.
When data are missing completely at random (MCAR), conventional boosting algorithms can be directly applied to the complete observations, yielding valid results as they still constitute a random subsample. However, directly applying boosting procedures to data missing not at random (MNAR) poses a significant challenge, as it can lead to biased results (Bian et al., 2025b). This issue is further complicated by the inherent non-identifiability problems associated with MNAR data, as discussed by Wang et al. (2014), Miao and Tchetgen Tchetgen (2016), Sun et al. (2018), and Morikawa and Kim (2021), among others.
In this paper, we incorporate two strategies to adjust the loss function: the Buckley-James-type and the inverse propensity weight adjustments, with the missingness effects accounted for. These adjustments allow the flexible application of boosting methods to incomplete data with MNAR through a functional gradient descent algorithm. However, the usual optimization procedures are complicated by the involvement of unknown functions. To address this difficulty, we describe semiparametric optimal estimation approaches that provide consistent estimators for these unknown functions. We rigorously establish the theoretical properties for the resultant estimators. In addition, we adapt conditions considered by Morikawa and Kim (2021) to ensure model identifiability.
@@ 2.1 Objective and Data
[…]
@@ 2.2 Conventional Boosting Prediction Procedure
Finding $\hat{f}_{\text{full}}$ in (3) can be achieved by employing the boosting method (Freund and Schapire, 1997). This method, also known as the functional gradient descent algorithm (e.g., Bühlmann and Yu, 2003; Bühlmann and Hothorn, 2007; Schapire and Freund, 2012), minimizes the empirical risk through steepest gradient descent in a function space to iteratively improve the estimate of $\hat{f}_{\text{full}}$ (Friedman, 2001), where the function space is spanned by a class of constrained continuous functions, denoted by $\mathcal{C}$. At iteration $m$, given the current estimate $f^{(m)}(\cdot)$, the next estimate $f^{(m+1)}$ of $\hat{f}_{\text{full}}$ is updated by adding an increment term, $\hat{\alpha}^{(m+1)}\hat{h}^{(m+1)}(\cdot)$, where $\hat{h}^{(m+1)}(\cdot)\in\mathcal{C}$, and $\hat{\alpha}^{(m+1)}$ is a learning rate.
@@
[…]
@@ 3.1 Missing not at Random and Identifiablity
Let $f(y|X=x)$ denote the conditional probability density of $Y$ given $X$, and let $\pi(y,x)\triangleq\Pr(R=1|Y=y,X=x)$ denote the propensity of observing the response, i.e., the conditional probability of observing $Y$, given $Y$ and $X$. If $\pi(y,x)$ does not depend on $Y$, the resulting missing data mechanism is called missing at random (MAR). When the missing mechanism is missing not at random (MNAR) or nonignorable, the propensity $\pi(y,x)$ depends on $Y$, as well as $X$, irrespective of whether $Y$ is missing or observed. Under MNAR, non-identifiability is often a concern (Robins and Ritov, 1997). When both $f(y|X=x)$ and $\pi(y,x)$ are left fully unspecified, the joint distribution of $Y$ and $R$, given $X$, becomes non-identifiable (Robins and Ritov, 1997).
To address non-identifiability issues, certain assumptions can be imposed. For example, Wang et al. (2014) assumed the existence of a nonresponse instrumental variable (aka a shadow variable) (Miao and Tchetgen Tchetgen, 2016), which is a component of the covariate vector $X$ that is associated with the response $Y$ but is conditional independent of the missingness indicator $R$, given $Y$ and other components of $X$. Alternatively, Sun et al. (2018) assumed the existence of another version of instrumental variables: a subset of the covariate vector $X$ that is independent of the response $Y$ but conditionally dependent on $R$, given $Y$ and other components of $X$. While utilizing instrumental variables can mitigate non-identifiability issues, identifying an appropriate instrumental variable can be difficult in applications. Most importantly, those conditions are not testable solely based on the observed data (Morikawa and Kim, 2021).
Instead of relying on the existence of instrumental variables, Morikawa and Kim (2021) proposed a set of identification conditions, listed as (A1) - (A4) in the Supplementary Materials. Those conditions are testable using observed data. Even if the model does not satisfy those conditions, a doubly-normalized exponential transformation (Morikawa and Kim, 2021) can be applied to artificially make the model identifiable, albeit at the cost of sacrificing the estimator consistency. In our development here, we adopt the identification conditions of Morikawa and Kim (2021) to emphasize our key ideas.
Imposing parametric assumptions on both $f(y|X=x)$ and $\pi(y,x)$ can help address the non-identifiability issue, but the results are often sensitive to model misspecification (Kenward, 1998). As a remedy, an intermediate approach may be considered, where one function is modeled parametrically while leaving the other unspecified; this is the strategy we adopt in the following development.
@@ 3.2 Adjusting Loss Functions with MNAR Data
The boosting prediction procedure described in Section 2.2 cannot be applied directly to $\mathcal{O}_{\text{missing}}$, as the response $Y_{i}$ is not observed for every subject in the study. To address this, we construct a new loss function, denoted $L^{*}(y_{i},f(x_{i}),r_{i})$, using the observed data in the sample $\mathcal{O}_{\text{missing}}$. Following the idea of Chen and Yi (2024), we construct an adjusted loss function $L^{*}(\cdot,\cdot,\cdot)$ that maintains the same risk function as the original loss function $L(\cdot,\cdot)$, i.e., $E\{L^{*}(Y_{i},f(X_{i}),R_{i})\}=E\{L(Y_{i},f(X_{i}))\}$. Therefore, minimizing the risk function of the adjusted loss function $E\{L^{*}(Y_{i},$ $f(X_{i}),R_{i})\}$ is equivalent to minimizing the risk function $R(f)$ defined in (1), as if all $Y_{i}$ were observed.
Following Hothorn et al. (2006), one way to adjust the loss function is through the inverse propensity weight (IPW) scheme:
@@
To maximize the use of all available measurements, we follow the idea of Chen and Yi (2024) to construct a Buckley-James (BJ)-type adjusted loss function:
@@
where $\Psi_{0}(x)\triangleq E\{L(Y,f(X))|X=x,R=0\}$, determined by $\int L(y,f(x))f(y|X=x,R=0)dy$.
Determining $\Psi_{0}(x)$ requires $f(y|X=x,R=0)$, which is not available since the outcome is not observed for this subpopulation. To get around this, as in Kim and Yu (2011), we use Bayes’ rule to express
@@ 4.1 Construction of Consistent Estimators
Since some $y_{i}$ values are missing (i.e., when $r_{i}=0$), directly setting (8) to 0 to solve for $\gamma$ is not feasible. To address this, Morikawa and Kim (2021) proposed an alternative estimation method. Let $S_{0}(Y,X;\gamma)$ denote (9) with $r$ set to 0. Let
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4.3 Specification of the class $\mathcal{C}$
The boosting prediction procedure for MNAR data described in Section 4.2 depends on the specification of the class $\mathcal{C}$. Similar to Chen and Yi (2024), we consider the regression spline method to characterize the functions in $\mathcal{C}$. Instead of using the truncated power basis functions, which are known to be numerically unstable in practice (Hastie et al., 2009), we employ more computationally stable B-spline basis functions (de Boor, 2001).
@@
with $h_{0}$ denoting the intercept and $h_{k}(x_{k})=\sum^{T}_{t=1}\xi_{kt}b_{kt}(x_{k})$ for $k=1,\ldots,p$, where $x=(x_{1},\ldots,x_{p})^{\top}$; $b_{k}(x_{k})=(b_{k1}(x_{k}),\ldots,$ $b_{kT}(x_{k}))^{\top}$ is the vector of the B-spline basis functions of order $M$, with TM+1T-M+1 interior knots at suitably chosen quantiles of $X_{k}$ and $M\geq 2$ being an integer; and $\xi_{k}=(\xi_{k1},\ldots,\xi_{kT})^{\top}$ is the unknown parameter vector.
@@ 5 Theoretical Results
While the convergence of Algorithm 1 is guaranteed by Theorem 1, in applications, Algorithm 1 stops at a finite number to avoid excessively running an unnecessarily large number of iterations. Similar to Chen and Yi (2024), we may use (18), in combination with (17), to decide an upper bound for the stopping iteration number $\tilde{m}$ in Algorithm 1:
@@
where $\eta$ is a given threshold in Algorithm 1. The upper bound of (19) shows that the number of iterations may vary with the choice of an initial function $f^{(0)}$ and the required accuracy $\eta$, as noted by Boyd and Vandenberghe (2004) and Chen and Yi (2024).
@@ 6.1 Simulation Design and Data Generation
We consider two settings, as in Bian et al. (2025b). with different dimensions $p$ for covariates $X_{i}$: Setting 1 has $p=2$, and Setting 2 includes a larger number of covariates with $p=9$.
@@
[…]
@@
Condition (A4) in the Supplementary Materials implies that setting $\beta_{1,3}\neq 0$ ensures the model to be identifiable. Noted by Morikawa and Kim (2021) through simulation studies, almost half of the simulations failed to converge if $\beta_{1,3}$ is taken as 0. Consequently, here we only consider settings with a nonzero $\beta_{1,3}$, set as $\beta_{1,3}=0.8$.